OpenStudy (anonymous):

HELP Please!!!!!!!! Ill Give Medal!!!!!! Solve for X x-8/2 = -3 A. 2 B. –14 C. –9 D. 7

OpenStudy (anonymous):

B. -14

OpenStudy (anonymous):

C. -9

OpenStudy (anonymous):

\[\frac{ x-8 }{ 2 } = -3\]

OpenStudy (anonymous):

So we need to isolate x by it's self right?

OpenStudy (anonymous):

if you give the answer explain how you got it atleast

OpenStudy (anonymous):

To get rid of that denominator of 2 we can multiply both sides by 2 so it cancels like

OpenStudy (anonymous):

ooh ok :)

OpenStudy (aaronandyson):

x-8/2 = -3 x-8 = 3 times 2 x-8 = -6 x = -6 + 8 x= 2

OpenStudy (anonymous):



OpenStudy (anonymous):

so x - 8 = -6

OpenStudy (anonymous):

:D thanks can u help with some more???

OpenStudy (aaronandyson):

@swagmaster47 i already explained it to her O.O

OpenStudy (anonymous):

I can

OpenStudy (aaronandyson):

Sure @Kayla369

OpenStudy (anonymous):

:D kk

OpenStudy (anonymous):

o n there is no -6

OpenStudy (anonymous):

That isnt your final answer

OpenStudy (anonymous):

x - 8 = -6

OpenStudy (anonymous):

you want x by it's self

OpenStudy (anonymous):

ah ok lol

OpenStudy (aaronandyson):

@Kayla369 x-8/2 = -3 x-8 = 3 times 2 x-8 = -6 x = -6 + 8 x= 2

OpenStudy (anonymous):

oh haha

OpenStudy (aaronandyson):

More Questions?

OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. y ÷ 2 x 4 - 10 ÷ x + 6 where x = 5 and y = 16 A. 36 B. 21 C. -3 D. 3

OpenStudy (aaronandyson):

?, i can only see "?" over there.

OpenStudy (anonymous):

y divided by 2 x 4 - 10 divided by x = 6 where x = 5 and y = 16

OpenStudy (anonymous):

ya ik i hate it when that happens

OpenStudy (aaronandyson):

I cant't understand the Question, take a screenshot of it and send it

OpenStudy (anonymous):

\[y \div 2 \times 4 - 10 \div x = 6 ,where, x = 5 ,and,y = 16\]

OpenStudy (anonymous):

one sec lol idk how to screen shot it so one sec

OpenStudy (aaronandyson):

substitute the values in

OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. \[y \div 2 \times 4 - 10 \div x + 6, where, x = 5 , and ,y = 16\] A. 36 B. 21 C. –3 D. 3

OpenStudy (anonymous):

lol nvm i just saw what u said :D

OpenStudy (aaronandyson):

Given y = 16 , x = 5 y divided by 2 times 4 minus 10 divided by x this mean 16 divided by 2 times 4 minus 10 divided 5

OpenStudy (anonymous):

Solve \[\frac{ z-5 }{ 7 } = -1\] z=????

OpenStudy (anonymous):

yep brb need to reload my page

OpenStudy (aaronandyson):

did you get the one you asked earlier?

OpenStudy (aaronandyson):

z-5/7 = -1 z-5 = -1 times 7 z-5 = -7 z = -7 + 5 z = -2